OpenStudy (anonymous):

Imagine the point (11 -8) on the terminal side of an angle. Find the value of the secant of this angle.

OpenStudy (amistre64):

just find the cosine and flip it over ...

OpenStudy (anonymous):

How do I do that?

OpenStudy (amistre64):

cosine is a basic trig function that is well defined in your material ... what are your 3 basic trig function definitions?

OpenStudy (anonymous):

Do you mean sin, cos, tan?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

lol, that actually got me thinking of something that might make this simpler to do without finding a missing side length

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

do you agree that tan(angle) = y/x ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

there is a pythagorean relationship that goes like this: \[sin^2+cos^2 = 1\] \[\frac{1}{cos^2}(sin^2+cos^2 = 1)\] \[tan^2+1 = sec^2\] \[\sqrt{tan^2+1} = sec\]

OpenStudy (amistre64):

we are given x and y to play with ... 11, -8.\[\sqrt{(\frac{-8}{11})^2+1}=sec\]

OpenStudy (amistre64):

or we could run thru finding the hypotenuse with:\[r=\sqrt{(-8)^2+(11)^2}\] \[r=\sqrt{64+121}\] \[r=\sqrt{185}\] and therefore cos = 11/sqrt(185) sec = sqrt(185)/11 simplify as wanted

OpenStudy (anonymous):

Thank you. That was so confusing .Thanks again.

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

I'll need it :)